# Multi-frequency analysis of the ALMA and VLA high resolution continuum observations of the substructured disc around CI Tau

### v1

@@ 2.1 Observations
We introduce, for the first time, ALMA Band 3, VLA Ka, Ku, X, and C Band observations of CI Tau. We make use of archival ALMA Band 6 (Konishi et al., 2018; Clarke et al., 2018) and 7 (Rosotti et al., 2021) observations to study the properties of solids in this disc at high angular resolution and sensitivity. Information about these datasets is summarised in footnote 2.
@@
The SB observations used an array with baseline lengths from $15.3\,{\rm m}$ to $2.4\,{\rm km}$ ($\approx 400\,{\rm mas}$ resolution) and 41 antennas. The total integration time on the science target was $\approx 38\,{\rm min}$. The phase calibrator J0426+2327 was observed in an alternating sequence with the science target, every $10\,{\rm min}$. The bandpass and flux calibrator J0237+2848 was observed at the beginning of the observing block. The LB observations used an array with baseline lengths from $83.1\,{\rm m}$ to $16.2\,{\rm km}$ ($\approx 50\,{\rm mas}$ resolution). Two execution blocks (EBs) were scheduled. The first used 42 antennas and the second 46. The total integration time on the science target was $\approx 44\,{\rm min}$ for each EB. The phase calibrator J0426+2327 was observed in an alternating sequence with the science target, every minute. The bandpass and flux calibrator, J0423–0120 and J0510+1800 for different EBs, respectively, was observed at the beginning of the observing block. An additional “check” calibrator, J0425+2235, was observed every $\approx 15\,{\rm min}$ to assess the quality of phase transfer.
@@
The publicly-available archival ALMA Band 6 observations (that we took into account) were conducted in Cycle 3 in August 2016 as part of the program 2015.1.01207.S (PI: H. Nomura) and in Cycle 4 in September 2017 as part of the program 2016.1.01370.S (PI: C. J. Clarke). The data were correlated from four SPWs in dual polarisation mode. For the program 2016.1.01370.S, all the SPWs were set in time division mode (TDM), each with 128 channels $15.625\,{\rm MHz}$ wide, spanning a total bandwidth of $2.0\,{\rm GHz}$, and centred at 224.0, 226.0, 240.0, and $242.0\,{\rm GHz}$. Instead, for the program 2015.1.01207.S, the SPWs were set in FDM and centred at: (i) $220.0\,{\rm GHz}$ with a bandwidth of $234.0\,{\rm MHz}$ in 480 channels of $488.3\,{\rm kHz}$; (ii–iii) 216.7 and $231.2\,{\rm GHz}$ with a bandwidth of $937.5\,{\rm MHz}$ in 1920 channels of $488.3\,{\rm kHz}$; (iv) $234.4\,{\rm GHz}$ with a bandwidth of $937.5\,{\rm MHz}$ in 960 channels of $976.6\,{\rm kHz}$.
@@
CI Tau was observed by the VLA in Ka Band in two configurations as part of the project 19A-440 (PI: M. Tazzari). The SPWs covered the frequency range between 29 and $37\,{\rm GHz}$ (wavelengths between 8.1 and $10.3\,{\rm mm}$). Observations in B configuration, with a maximum baseline of $11.1\,{\rm km}$, were performed during March/April 2019 and September/October 2020, for a total on-source integration time of $9.8\,{\rm hrs}$. Observations in A configuration, with a maximum baseline of $36.6\,{\rm km}$, were executed between September and October 2019, resulting in a total on-source integration time of $8.6\,{\rm hrs}$. For both configurations, 3C147 served as the flux and bandpass calibrator, J0403+2600 was used for pointing calibration, and J0431+2037 for phase calibration.

### v2

@@ 2.1 Observations
We introduce, for the first time, ALMA Band 3, VLA Ka, Ku, X, and C band observations of CI Tau. We make use of archival ALMA Band 6 (Konishi et al., 2018; Clarke et al., 2018) and 7 (Rosotti et al., 2021) observations to study the properties of solids in this disc at high angular resolution and sensitivity. Information about these datasets is summarised in footnote 21.
@@
The SB observations used an array with baseline lengths from $15.3\,{\rm m}$ to $2.4\,{\rm km}$ ($\approx 400\,{\rm mas}$ resolution) and 41 antennas. The total integration time on the science target was $\approx 38\,{\rm min}$. The phase calibrator J0426+2327 was observed in an alternating sequence with the science target, every $10\,{\rm min}$. The bandpass and flux calibrator J0237+2848 was observed at the beginning of the observing block. The LB observations used an array with baseline lengths from $83.1\,{\rm m}$ to $16.2\,{\rm km}$ ($\approx 50\,{\rm mas}$ resolution). Two execution blocks (EBs) were scheduled. The first used 42 antennas and the second 46. The total integration time on the science target was $\approx 44\,{\rm min}$ for each EB. The phase calibrator J0426+2327 was observed in an alternating sequence with the science target, every minute. The bandpass and flux calibrator, J0423–0120 and J0510+1800 for different EBs, respectively, was observed at the beginning of the observing block. An additional ‘check’ calibrator, J0425+2235, was observed every $\approx 15\,{\rm min}$ to assess the quality of phase transfer.
@@
The publicly available archival ALMA Band 6 observations (that we took into account) were conducted in Cycle 3 in August 2016 as part of the program 2015.1.01207.S (PI: H. Nomura) and in Cycle 4 in September 2017 as part of the program 2016.1.01370.S (PI: C. J. Clarke). The data were correlated from four SPWs in dual polarisation mode. For the program 2016.1.01370.S, all the SPWs were set in time division mode (TDM), each with 128 channels $15.625\,{\rm MHz}$ wide, spanning a total bandwidth of $2.0\,{\rm GHz}$, and centred at 224.0, 226.0, 240.0, and $242.0\,{\rm GHz}$. Instead, for the program 2015.1.01207.S, the SPWs were set in FDM and centred at: (i) $220.0\,{\rm GHz}$ with a bandwidth of $234.0\,{\rm MHz}$ in 480 channels of $488.3\,{\rm kHz}$; (ii–iii) 216.7 and $231.2\,{\rm GHz}$ with a bandwidth of $937.5\,{\rm MHz}$ in 1920 channels of $488.3\,{\rm kHz}$; (iv) $234.4\,{\rm GHz}$ with a bandwidth of $937.5\,{\rm MHz}$ in 960 channels of $976.6\,{\rm kHz}$.
@@
CI Tau was observed by the VLA in Ka band in two configurations as part of the project 19A-440 (PI: M. Tazzari). The SPWs covered the frequency range between 29 and $37\,{\rm GHz}$ (wavelengths between 8.1 and $10.3\,{\rm mm}$). Observations in B configuration, with a maximum baseline of $11.1\,{\rm km}$, were performed during March and April 2019 and September and October 2020, for a total on-source integration time of $9.8\,{\rm hrs}$. Observations in A configuration, with a maximum baseline of $36.6\,{\rm km}$, were executed between September and October 2019, resulting in a total on-source integration time of $8.6\,{\rm hrs}$. For both configurations, 3C147 served as the flux and bandpass calibrator, J0403+2600 was used for pointing calibration, and J0431+2037 for phase calibration.